I need help solving this problem? y(9+y)=4(2y+5)
Answer by Mr. Math Wiz(7)

-First, you have to factor the numbers outside the parenthesis into the parenthesis
y(9)=9y, y(y)=y^2, 4(2y)=8y, 4(5)=20
-So you get (9y+y^2)=(8y+20), now you have to figure out y so subtract the 8y from the 9y
-9y+y^2=8y+20
-8y -8y
y+y^2=20
-Move the twenty to the other side by subtraction
y^2+y-20=0 <----ANSWER
HOPE THIS HELPED :)